OpenStudy (anonymous):

Help please? What is the square root of 54x7?

OpenStudy (anonymous):

Hey friend ! First we have 54 = 2 . 27 and 54.7 means : 2.27.7 and (2.27.7)^2 too ! Now answer !

OpenStudy (debbieg):

7 is already prime. Fully factor 54, and you should easily see what (if any) factors of it are perfect squares. Remember:\[\Large \sqrt{a\cdot b }=\sqrt{a }\cdot \sqrt{ b }\] So once you fully factor the stuff under the radical sign, just "pull out" the square roots of anything that is a perfect square, and "leave alone" anything that isn't (it has to stay under the radical).